OpenStudy (ravneet):

tan 10 - tan 50 +tan 70 with trigonometry answer is sqrt(3) but i cant solve it

OpenStudy (anonymous):

\[\tan 10 - \tan 50 +\tan 70 = \sqrt{3}\] IS THIS WHAT YOUR SAYING?

OpenStudy (ravneet):

ya

OpenStudy (raden):

tan70 = tan(60+10) use the identity : tan(a+b) = (tan(a) + tan(b))/(1-tan(a)tan(b)) so, tan70 = (tan60+tan10)/(1-tan60tan10) tan70 = (sqrt(3)+tan(10))/(1-sqrt(3)tan10) ... (1) tan50 = tan(60-10) use the identity : tan(a-b) = (tan(a) - tan(b))/(1+tan(a)tan(b)) so, tan50 = (tan60-tan10)/(1+tan60tan10) tan50 = (sqrt(3)-tan(10))/(1+sqrt(3)tan10) ... (2) therefore, the value of tan70 - tan50 = (sqrt(3)+tan(10))/(1-sqrt(3)tan10) - (sqrt(3)+tan(10))/(1-sqrt(3)tan10) if that simplied becomes, we get tan70 - tan50 = 8tan10/(1-3tan^2 10)

OpenStudy (raden):

now, look at the original question above is , sin10-sin50+tan70 which similar with tan70 - tan50 + tan10. (the value of tan70-tan50 already done) so, tan70 - tan50 + tan10 = 8tan10/(1-3tan^2 10) + tan10 = (8tan10+tan10-3tan^3 10)/(1-3tan^2 10) = (9tan10-3tan^3 10)/(1-3tan^2 10) = 3 (3tan10 - tan^3 10)/(1-3tan^2 10)

OpenStudy (raden):

then, remember that there is a formula in trigono : tan(3x) = (3tan(x) - tan^3 (x))/(1 - 3tan^2 (x))

OpenStudy (raden):

therefore, the calculation above can be simplied becomes 3 (3tan10 - tan^3 10)/(1-3tan^2 10) = 3 tan(3*10) = 3 tan30 = 3 * 1/3 sqrt(3) = sqrt(3) QED ^_^